given the function f described by f(x)=x+18 find each of the following
f(0)= 18
f(-15)=-15+18=3
f(-14)=-14+18=4
f(5)=5+18=23
f(c+12)=(c+12)+18= c+30

Answer by curioustoknow(5)

here is the answers:
f(x)=x+18 find each of the following
f(0)= 18
f(-15)=-15+18=3
f(-14)=-14+18=4
f(5)=5+18=23
f(c+12)=(c+12)+18= c+30